How much water should be added to 100 liters of 90% alcohol concentration to dilute it to 75% alcohol concentration?
Answer by Maths68(1474)

Solution A
Amount = x
Concentration =0% =0 (It is pure water so concentration of Acohoal wil be zero)
Solution B
Amount = 100
Concentration =90% = 0.9
Resultant Solution
Amount =100+x
Concentration =75%=0.75
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
(0)x+(0.9)(100)=(100+x)(0.75)
90=75+0.75x
90-75=0.75x
15=0.75x
15/0.075=x
20=x
Solution A
Amount = x =20 liters
Concentration =0% =0 (It is pure water so concentration of Acohoal wil be zero)
Solution B
Amount = 100
Concentration =90% = 0.9
Resultant Solution
Amount =100+x =100+20=120 liters
Concentration =75%=0.75

Result
---------
20 liters water should be added to 100 liters of 90% alcohol concentration to dilute it to 75% alcohol concentration.